# A deep dive into the interplay of structured quantum peaked circuits and infinite temperature correlation functions

### v1

@@ Abstract
Random quantum circuits have been extensively explored for quantum supremacy demonstrations [1, 2, 3, 4]. However, verifying their output distributions remains challenging [2, 5, 6, 7]. Here, we propose the infinite-temperature correlation function (ITCF) as a physically meaningful observable for noisy intermediate-scale quantum (NISQ) devices one that can be extracted using engineered circuits rather than relying on fully random constructions. This is realized by leveraging peaked quantum states whose probability distributions are sharply peaked at specific outcomes due to constructive interference thus offering more efficient verifiability and stronger signal observability. Rather than using Haar-random states, which often yield vanishing signals through destructive interference, we construct purposefully biased quantum states using either Grover-based amplitude amplification or shallow structured circuits. These engineered states amplify contributions from relevant operator subspaces, enabling robust detection of non-zero ITCF values that would otherwise be suppressed under random-state sampling. Our results highlight a problem-specific state preparation framework that mitigates signal loss from random averaging and facilitates the detection of physically meaningful observables in NISQ devices. We also discuss future extensions to multi-qubit observables, scrambling diagnostics, and variational circuit optimization, underscoring the broader potential of Peaked States for quantum simulation and verification.
@@ I Introduction
Random quantum circuits have played a central role in recent efforts to demonstrate quantum advantage, particularly in the context of sampling tasks believed to be beyond the reach of classical computers[1, 3, 4]. By drawing from the vast space of random unitary operations, these circuits generate highly entangled states that are exceedingly difficult for classical hardware to simulate. However, verifying these outputs—essential for both complexity-theoretic claims and practical certification—remains a formidable hurdle, especially as the system size grows, due to the inherent randomness and lack of physically interpretable structure in the generated states [2, 5, 6, 7]. Recent work further reveals that the output distributions of RCS circuits undergo distinct phase transitions as a function of circuit depth and noise levels, thereby reinforcing the classical intractability of simulating such circuits even under moderate noise conditions [8].
@@
This verification challenge prompts a broader question: how can we extract physically meaningful observables from near-term, noisy intermediate-scale quantum (NISQ) devices when random sampling schemes often yield exponentially small signals? One promising response comes from the idea of peaked circuits—quantum circuits that produce probability distributions sharply biased toward specific bitstrings. Originally proposed to improve the verifiability of random circuit sampling [9], such circuits boost the likelihood of “heavy” outputs (i.e., bit-strings that occur with above-median probability) in a complexity-theoretic context. Crucially, the underlying principle—biasing quantum states to emphasize certain outcomes—extends beyond complexity theory and suggests a broader class of applications. In particular, it motivates the exploration of quantum observables that are both sensitive to a system’s dynamics and practically measurable using structured, non-random circuits.
@@
One such observable is the infinite-temperature correlation function (ITCF), which captures the dynamical relationship between quantum operators at different times under thermal equilibrium at infinite temperature. ITCFs have been widely employed in studies of quantum chaos, thermalization, and information scrambling [10, 11], offering a conceptually clear and hardware-compatible tool for probing many-body quantum dynamics. In this work, we specifically focus on the equal-time case ($t=0$), as it presents an experimentally accessible scenario on current NISQ hardware while still capturing essential aspects of operator correlations. However, when estimated using Haar-random or uniformly sampled states, ITCFs often suffer from destructive interference and signal cancellation across the Hilbert space, leading to vanishing expectation values that are difficult to detect on today’s quantum devices.
To overcome this challenge, we introduce the notion of peaked quantum states—quantum states engineered to amplify specific outcomes through constructive interference. Rather than relying on random-state preparations, we generate these states using structured quantum circuits designed to enhance contributions from operator-relevant subspaces. We explore two complementary strategies: (1) Grover-based amplitude amplification, which systematically boosts amplitudes associated with selected computational basis states, and (2) shallow, engineered circuits that preferentially peak probability distributions toward physically meaningful components. These peaked states enable direct measurement of non-vanishing ITCF signals with significantly improved signal-to-noise ratios, even in the presence of noise and imperfections characteristic of NISQ devices. Our approach also involves a mathematical reformulation of the ITCF estimator that allows peaked quantum states to efficiently target operator-relevant contributions within the Hilbert space—bridging the gap between circuit design and physical observability. In contrast to prior work relying on deep or variationally optimized circuits [9], our method offers a lightweight and interpretable alternative. Our experiments on IBM Q devices confirm that structured peaked circuits can robustly amplify ITCF signals, providing a practical demonstration of their applicability in realistic noisy environments.
In what follows, we describe the construction of such circuits, analyze their behavior theoretically and experimentally, and discuss extensions to multi-qubit observables, scrambling diagnostics (e.g., out-of-time-order correlators), and hybrid variational protocols that could further enhance peaked-state design.
@@ II-B Reformulation of the Estimator
Here, $P_{\uparrow}$ is the projector onto the $+1$ eigenstate of the Pauli-$Z$ operator, satisfying $P_{\uparrow}^{2}=P_{\uparrow}$. This leads to the following reformulated estimator:
@@
Hence, convergence improves with system size, and in some cases, a single sample suffices. The estimator can then be rewritten as:
@@
By the same logic, Equation (10) can also be manipulated given the structure $A(t)=V(-t)A(0)V(t)$. If we again expand terms in $V(t)$ up to zeroth order in time then the quantity transforms to $\langle r|A(0)|r\rangle$. We leave it as that without expanding it further in the computational basis. Consequently, the full estimator can be approximated as:
@@ II-C Limitations of Haar Sampling
  • Limited shots on NISQ devices: In practice, measurement repetitions are constrained, and statistical convergence becomes difficult [5].
@@ III-B Grover-Based Circuit: Structured Peaking
Output: Peaked quantum state |$\ket{r}$
1:Initialize all qubits to |$\ket{0}^{\otimes n}$
@@ III-C Shallow Circuit: Lightweight Peaking
Output: Peaked quantum state |$\ket{r}$
1:Initialize all qubits to |$\ket{0}^{\otimes n}$
@@
    Partial Superposition: Apply Hadamard only to a subset of qubits that significantly affect the observable $A(0)$, leaving the others in the |$\ket{0}$ state. For example, suppose the observable is $A(0)=Z_{0}\otimes I_{1}\otimes Z_{2}\otimes I_{3}\otimes Z_{4}$. Then the expectation value $\langle z_{j}|A(0)|z_{j}\rangle$ depends only on qubits 0, 2, and 4. In this case, Hadamard gates are applied only to those three qubits to create superposition, while qubits 1 and 3 are left unentangled in |$\ket{0}$. This reduces circuit depth and avoids introducing noise from qubits that do not influence the measurement.
@@
    Targeted Bias: For qubits expected to contribute significantly to the value of $\langle z_{j}|A(0)|z_{j}\rangle$, apply $R_{y}(\theta_{i})$ rotations to tilt their amplitudes toward |$\ket{0}$ or |$\ket{1}$. Continuing the above example, states with even parity on qubits 0, 2, and 4 (i.e., $z_{0}+z_{2}+z_{4}$ is even) lead to a positive expectation value. To amplify such configurations, $R_{y}(\pi/4)$ rotations are applied to qubits 0, 2, and 4. This biases the probability distribution toward favorable basis states and enhances the signal-to-noise ratio in downstream correlation estimates.

### v2

@@ Abstract
Random quantum circuits have been extensively explored for quantum supremacy demonstrations [1, 2, 3, 4]. However, verifying their output distributions remains challenging [2, 5, 6]. Here, we propose the infinite-temperature correlation function (ITCF) as a physically meaningful observable for noisy intermediate-scale quantum (NISQ) devices one that can be extracted using engineered circuits rather than relying on fully random constructions. This is realized by leveraging peaked quantum states whose probability distributions are sharply peaked at specific outcomes due to constructive interference thus offering more efficient verifiability and stronger signal observability. Rather than using Haar-random states, which often yield vanishing signals through destructive interference, we construct purposefully biased quantum states using either Grover-based amplitude amplification or shallow structured circuits. These engineered states amplify contributions from relevant operator subspaces, enabling robust detection of non-zero ITCF values that would otherwise be suppressed under random-state sampling. Our results highlight a problem-specific state preparation framework that mitigates signal loss from random averaging and facilitates the detection of physically meaningful observables in NISQ devices. We also discuss future extensions to multi-qubit observables, scrambling diagnostics, and variational circuit optimization, underscoring the broader potential of Peaked States for quantum simulation and verification.
@@ I Introduction
Random quantum circuits have played a central role in recent efforts to demonstrate quantum advantage, particularly in the context of sampling tasks believed to be beyond the reach of classical computers [1, 3, 4]. By drawing from the vast space of random unitary operations, these circuits generate highly entangled states that are exceedingly difficult for classical hardware to simulate. However, verifying these outputs—essential for both complexity-theoretic claims and practical certification—remains a formidable hurdle, especially as the system size grows, due to the inherent randomness and lack of physically interpretable structure in the generated states [2, 5, 6]. Moreover, recent studies show that output distributions undergo sharp phase transitions with depth and noise, further reinforcing their classical inaccessibility [7].
@@
This verification challenge raises a broader question: how can we extract physically meaningful observables from near-term, noisy intermediate-scale quantum (NISQ) devices, when random sampling often yields exponentially small signals? A promising direction is the use of peaked circuits—quantum circuits that bias probability distributions toward specific bitstrings. Initially introduced to improve the verifiability of random circuit sampling [8], they boost the likelihood of “heavy” outputs in a complexity-theoretic sense. More broadly, this biasing principle motivates new applications: designing circuits that amplify physically informative components of quantum observables in structured and experimentally feasible ways.
@@
One such observable is the infinite-temperature correlation function (ITCF), which captures the dynamical relationship between quantum operators at different times under thermal equilibrium at infinite temperature. ITCFs have been widely employed in studies of quantum chaos, thermalization, and information scrambling [9, 10], offering a conceptually clear and hardware-compatible tool for probing many-body quantum dynamics. In this work, we specifically focus on the equal-time case ($t=0$), as it presents an experimentally accessible scenario on current NISQ hardware while still capturing essential aspects of operator correlations. However, estimating ITCFs using Haar-random or uniformly sampled pure states often leads to destructive interference across the Hilbert space, causing the expectation value to vanish exponentially with system size—even under ideal, noiseless conditions. This makes the full-space trace definition of ITCF, while theoretically well-defined, practically inaccessible in realistic experiments where only a limited number of samples can be taken.
To overcome this challenge, we propose a practical estimator that selectively amplifies operator-relevant contributions, thereby avoiding the need to average over the full Hilbert space. This is achieved using peaked quantum states—engineered to concentrate probability on physically informative basis components while preserving the core structure of the ITCF. We explore two complementary strategies: (1) Grover-based amplitude amplification, and (2) shallow, structured circuits tailored to emphasize operator-relevant basis states. While our estimator does not reproduce the exact, unbiased trace over the full Hilbert space, it provides a biased but operationally useful alternative that captures operator-relevant dynamics under realistic constraints. This tradeoff—between theoretical generality and experimental feasibility—defines the scope of our contribution. In contrast to prior work relying on deep or variationally optimized circuits [8], our method offers a lightweight and interpretable approach. Experiments on IBM Q devices confirm that structured peaked circuits can robustly amplify relevant correlation signals, offering a practical pathway to probe many-body dynamics in today’s noisy quantum environments.
@@ II-B Reformulation of the Estimator
Here, $P_{\uparrow}$ denotes a single-qubit projector onto the +1+1 eigenstate of the Pauli-$Z$ operator, i.e., $|0\rangle\langle 0|$, acting on a qubit that is distinct from those on which $A(t)$ acts. This separation ensures that the projector selects a meaningful subspace without trivially commuting with the observable. It satisfies $P_{\uparrow}^{2}=P_{\uparrow}$, and leads to the following reformulated estimator:
@@
Hence, convergence improves with system size, and in some cases—particularly when variance is sufficiently suppressed due to concentration of measure—a single sample may suffice. This behavior is similar to that of thermal pure quantum (TPQ) states [12], though our peaked states are not formally TPQ, but are engineered to amplify operator contributions in ITCF estimation. Therefore, when the system is large enough, the estimator simplifies to:
@@
By the same logic, Equation (10) can also be manipulated given the structure A(t)=V(t)A(0)V(t)A(t)=V(-t)A(0)V(t). If we again expand terms in $V(t)$ up to zeroth order in time then the quantity transforms to $\langle r|A(0)|r\rangle$. We leave it as that without expanding it further in the computational basis. Consequently, the full estimator can be approximated as:
@@ II-C Limitations of Haar Sampling
  • Limited shots on NISQ devices: In practice, measurement repetitions are constrained, and statistical convergence becomes difficult.
@@ III-B Grover-Based Circuit: Structured Peaking
[…]
@@ III-C Shallow Circuit: Lightweight Peaking
[…]
@@
[…]
@@
[…]